OpenStudy (anonymous):

Find the vertices, foci, and the equations of the asymptotes for the hyperbola.. INCOMING EQUATION

OpenStudy (anonymous):

\[x^{2}\div36 - y^{2}\div25 = 1\]

OpenStudy (anonymous):

So the foci. That's a tricky one. Why are there more than one?

OpenStudy (dumbcow):

like the name....and the song ok since the x term is positive the hyperbola opens left/right a = 6 b = 5 center: (0,0) vertices : (+-6 , 0) to find foci, solve for c a^2 + b^2 = c^2 36 +25 = c^2 c = sqrt61 foci: (+-sqrt61 , 0) asymptotes: y = +- (b/a)x y = +-(5/6)x

OpenStudy (dumbcow):

there are 2 foci because a hyperbola has 2 parabolas opening in opposite directions thus 2 vertices and 2 foci

OpenStudy (anonymous):

The vertice is 6 because square root of 36 is 6 right? That's how you determine that?

OpenStudy (dumbcow):

yes , the numbers underneath represent a^2 and b^2

OpenStudy (anonymous):

so why isn't it (6,5)

OpenStudy (dumbcow):

great question, a is always associated with the x_term here is a reference http://tutorial.math.lamar.edu/Classes/Alg/Hyperbolas.aspx

OpenStudy (dumbcow):

just remember : a goes with x determines horizontal vertices/foci b goes with y determines vertical vertices/foci